Jojobrowm:

sarai went an amusement park, paying $5 to get in. After that you paid $3 per ride. write the equation in slope intercept form to represent the situtation

kaneki:

how many rides were there?

jhonyy9:

y = mx +b m = 3 b = 5 hope this help you @Laylalyssa

supie:

So slope intercept form formula is \(y=mx+b\) \(m=3\\b=5\) PLug in `m`(slope) and `b` What you would have to do is plug in `3` for `m` and `5` for `b` So basically substitute 3 for m and 5 for b in the formula \(y=mx+b\) So then you get your equation what I would do is graph the equation then find the `y-intercept` `y-intercept` is where the line that was graphed intercepts with the y-axis

supie:

@jhonyy9

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie So slope intercept form formula is \(y=mx+b\) \(m=3\\b=5\) PLug in `m`(slope) and `b` What you would have to do is plug in `3` for `m` and `5` for `b` So basically substitute 3 for m and 5 for b in the formula \(y=mx+b\) So then you get your equation what I would do is graph the equation then find the `y-intercept` `y-intercept` is where the line that was graphed intercepts with the y-axis \(\color{#0cbb34}{\text{End of Quote}}\) congrats !!! like a math teacher - very very nice !!!

supie:

Thanks!!